Steven Company has fixed costs of $276,000. The unit selling price, variable cost per unit, and contribution margin per unit for the company's two products are provided below.

Product Selling Price per Unit Variable Cost per Unit Contribution Margin per Unit
X $864 $324 $540
Y 731 391 340

The sales mix for product X and Y is 60% and 40% respectively. Determine the break-even point in units of X and Y combined. Round answer to nearest whole number.
units
Answer #1

Answer:

Unit selling price of sale mix = (864*60%)+(731*40%)=811

Unit variable cost of sales mix =(324*60%)+(391*40%)=351

Unit contribution Margin of sales mix = (811-351)= 460

Break Even sales unit = (276000/460)=600

600*60%=360 unit of X

600*40%= 240 unit of Y